OpenStudy (anonymous):

A construction worker plans to build a diagonal walking path through a rectangular park. If the dimensions of the park are 10 yards × 14 yards, what will the length of the path be? a. Draw a right triangle to represent this situation. Label the lengths of the known sides. Label the length of the unknown side x. b. Use the Pythagorean theorem to find the length of the missing side. Be sure to indicate which root represents the answer and write the answer in simplest radical form.

OpenStudy (anonymous):

Use a calculator to change the answer to decimal form. Round the answer to the nearest tenth. Using this decimal number, write a sentence to explain the meaning of the answer in this situation.

OpenStudy (anonymous):

@aum

OpenStudy (nikato):

a. Draw a right triangle where one leg is 10 and the other is 14. Label the hypothenuse(the diagonal) x becuz it's unknown

OpenStudy (anonymous):

oh okay

OpenStudy (nikato):

And pythagoreom theorem is (Leg1)^2 + (leg2)^2 = (hypothenuse)^2 Plug in your numbers (10)^2 + (14)^2 =x^2